OpenStudy (anonymous):

Which symbol makes the following TRUE? 6-4(3)______5-10 A.< B.> C.= D.none of the above Help meee >.<

OpenStudy (anonymous):

wheres the equation?

OpenStudy (anonymous):

oh, ok plug in each answer and solve until one is correct

OpenStudy (anonymous):

Kay Thank You !

OpenStudy (anonymous):

6-4(3)_(<)_5-10 6-12<-5 -6<-5 like that

OpenStudy (anonymous):

6-4(3)=-6 5-10=-5 -6<-5

OpenStudy (anonymous):

@B.Sanders why did you change the sign..?

OpenStudy (anonymous):

I didn't I was showing that 6-4(3) equaled -6 and 5-10 equaled -5 and then that -6 was less then -5?